Book says to solve: 1 over x = 2 over 3-x. I know you multiply (-1) into 2 over 3 - x but I'm lost after that. Can you help me?
Thanks

Answer by jorel1380(3719)

1/x=2/(3-x)
3-x=2x
3=3x
x=1..